# A comparative high-resolution spectroscopic analysis of in situ and accreted globular clusters

### v1

@@ 3.2 Deriving the abundances
The determination of chemical abundances of Mg, Si, Ca, TiI, TiII, Cr, Fe, Ni, and Zn was conducted by means of a comparative analysis between the observed equivalent widths (EWs), measured with the code DAOSPEC (Stetson & Pancino 2008) through the automatic tool 4DAO (Mucciarelli 2013), and theoretical line strengths. This analysis was executed using the code GALA (Mucciarelli et al. 2013).
We employed the spectral synthesis using the proprietary code SALVADOR to derive the chemical abundances for the species that have hyperfine/isotopic splitting transitions (ScII, V, Mn, Co, Cu, YII, BaII, LaII, and EuII). SALVADOR runs a -minimization between the observed line and a grid of suitable synthetic spectra computed on-the-fly by the code SYNTHE (Kurucz 2005) varying only the abundance of the matching element and keeping the stellar parameters fixed. Synthetic spectra were computed using all the atomic and molecular transitions available in the Kurucz/Castelli222http://wwwuser.oats.inaf.it/castelli/linelists.html database.
We used ATLAS9 (Kurucz 2005) model atmospheres computed assuming plane-parallel geometry, hydrostatic and radiative equilibrium, and local thermodynamic equilibrium for all the chemical elements. We started from an -enhanced model atmosphere ([/Fe] = +0.4 dex) for all the target GCs except for Rup106 for which we used a solar-scaled chemical mixture, according to the values of we derived. Examples of fits around absorption lines of elements of interest (i.e. Zn and EuII, see Sections 4.2 and 4.3) are displayed in Figure 2. Different synthetic spectra are computed for each star by assuming the atmospheric parameters derived as described in Section 3.1, and varying the abundances of Zn and EuII.
@@ 3.3 Abundance uncertainties
The uncertainties attributed to EW measurements have been estimated as the dispersion observed around the mean of individual line measurements, divided by the root mean square of the number of lines employed in the analysis.
The internal error on chemical abundances derived with spectral synthesis have been quantified using Monte Carlo simulations. To do so, we repeated the analysis on 1000 noisy synthetic spectra obtained adding Poissonian noise in order to mimic the S/N of the real spectra. The uncertainty is estimated as the standard deviation of the abundance distribution of the 1000 noisy synthetic spectra.
Finally, we performed new calculations of the chemical abundances, considering also the uncertainties associated with the atmospheric parameters. This involved varying one stellar parameter at a time while keeping the others fixed. In the end, all the sources of error are added in quadrature. Given that our chemical abundances are expressed as abundance ratios ([X/Fe]), the uncertainties on the iron abundance have been taken into account.
The total uncertainty has been estimated as the squared sum of these two components. Therefore, the final errors have been calculated as:
where is the dispersion around the mean of chemical abundances, is the number of used lines and are the abundance differences obtained by varying the parameter .
@@ 4 Abundances of target GCs
In this section, we discuss abundances of -, iron-peak, and neutron-capture elements of target GCs. We exercise caution when interpreting the abundances of certain chemical elements that may be influenced by the presence of multiple populations within GCs (e.g., O, Na, Mg, and Al, see Bastian & Lardo 2018). Among these species, we selectively focus only on Mg, which exhibits minimal dispersion within five out of six GC in our sample. This outcome aligns with expectations, particularly considering that targeted systems belong to the low-mass and high metallicity regime of the MW GC system, where the efficiency of the MgAl burning channel diminishes (Ventura et al. 2013, Dell’Agli et al. 2018, Alvarez Garay et al. 2024). Indeed, we find spread in the [Mg/Fe] distribution of stars only in the most massive among our target cluster, NGC 6626. Since we only care about the pristine chemical composition of the gas that formed the cluster, we derived the mean value of [Mg/Fe] only selecting first generation stars. The results obtained from the homogeneous chemical analysis of 66 RGB stars of the sample of six GCs are listed in Tables 2 - 4. Mean abundances for each GC are available in Table 5. The average abundance ratios of the target GCs are illustrated in Figures 3 - 5.
@@
In Figure 3 we show the mean values of the -elements abundances. These chemical elements are predominantly synthesized in massive stars and subsequently dispersed into the interstellar medium through core-collapse supernovae events (CC-SNe, Kobayashi & Nomoto 2009, Romano et al. 2010, Kobayashi et al. 2020). Additionally, there is a minor but discernible contribution from Type Ia supernovae (SNe Ia), particularly notable for Ca and Ti, as discussed in Kobayashi et al. (2020).
First of all, we find that the two GCs associated to GSE (green filled symbols) show identical abundances, within the uncertainties, in all the analysed -elements. Second, we show that mean abundances of Mg, Si and Ca of the three in situ GCs NGC 6218, NGC 6522 and NGC 6626 (red, brown and pink filled symbols, respectively) are fully compatible among each other. Ti is the element where such a consistency is less evident. Third, we find that NGC 362 and NGC 1261 are -depleted compared to the in situ GCs. They display differences in the -element at the level in Mg and Ca when comparing the average abundances and the respective standard deviations (see Appendix A.1 for a comparison with literature). Last, we derive subsolar abundances in the -elements of Rup106 (blue filled squares), specifically with , in good agreement with literature works (Brown et al. 1997, V13, see Appendix A.2 for a comparison with the results by V13). This is, on average, lower than the other five target clusters.
These results are particularly relevant in understanding to which degree can we use -elements to distinguish the origin of GCs. The observed trend matches well the expectations given by the cluster chrono-dynamical associations. Indeed, GSE clusters NGC 1261 and NGC 362 share the same -element abundances, as expected for clusters accreted from the same progenitor, with lower values compared to NGC 6218, NGC 6522 and NGC 6626. In fact, the in situ GCs consistently show higher abundances, as predicted for systems born in the MW at this metallicity, where the contribution of SNIa to the gas chemical enrichment was not yet significant. Finally, Rup106 has even lower compared to the other target GCs, differentiating from the trends displayed by GSE GCs, indicating an even more different birth environment, probably characterised by a very low star formation efficiency.
@@ 4.2 Iron-peak elements
Iron-peak elements primarily originate in Type II CC-SNe and hypernovae (HNe), with some contribution also from SNe Ia. Specifically, elements like Sc, Cu, and Zn are predominantly synthesized by massive stars, while for V and Co the contribution of SNe Ia is not negligible. On the other hand, Cr, Mn, and Ni are primarily produced by SNe Ia (Romano et al. 2010, Kobayashi et al. 2020).
The results shown in Figure 4 indicate that most of the iron-peak elements do not provide effective and clear discrimination between in situ and GSE clusters. Indeed, target GCs display coherent values in all the chemical spaces with the exception once again of Rup106, which is underabundant in ScII, V, Mn, Co, and Ni. In the bottom panel of Figure 4 we highlight the two elements where Rup106 is most depleted compared to the other systems. Specifically, Cu is the most underabundant among the chemical elements derived in this work ([Cu/Fe] for Rup106). Such low values of [Cu/Fe] can be reproduced in stellar systems with chemical evolutionary model assuming very inefficient star formation (star formation rate , Mucciarelli et al. 2021b). Thus, the pronounced depletion in all of these elements might indicate that Rup106 was born in an environment where the contribution of massive stars was extremely low.
Among the iron-peak elements, the only one demonstrating a statistically significant difference between accreted and in situ GCs is Zn: all the in-situ GCs are overabundant with respect to GSE GCs and Rup106 at the 1 level. In passing, we note that NGC 6626 shows a significant spread in [Zn/Fe], as already found in Villanova et al. (2017). Zn seems therefore a good tracer of the in situ or accreted origin of GCs, but it is not sensitive enough to help distinguishing among different accreted progenitors. Such a conclusion has been drawn also by Minelli et al. (2021), even though they use a sample of GCs at higher metallicity. Indeed, high [Zn/Fe] ratios are expected for in situ GCs, and the rationale lies in the fact that the primary contributors to Zn production are hypernovae, which are linked to stars with masses around (Romano et al. 2010, Kobayashi et al. 2020). Consequently, galaxies with lower star formation rate should manifest lower [Zn/Fe] ratios due to the reduced impact of massive stars (Yan et al. 2020).
@@ 4.3 Neutron-capture elements
Slow (-) neutron-capture elements are mainly synthesized in Asymptotic Giant Branch (AGB) stars. In particular, light -process elements, (e.g., Y), are produced to a large extent by intermediate-mass AGB stars. Conversely, heavy -process elements (e.g., Ba, La) are mainly produced in AGB stars with masses lower than (Kobayashi et al. 2020). Rapid (-) neutron-capture elements (e.g., Eu), are synthesized during a broad spectrum of events, like special types of CC-SNe (rCCSNe, e.g., magnetorotational SNe, collapsars, Mösta et al. 2018, Siegel et al. 2019, Kobayashi et al. 2020) and neutron-star mergers (NSM, Lattimer & Schramm 1974).
@@
First of all, we observe that the in situ NGC 6218 is enhanced () in [YII/Fe] compared to the other two in situ GCs and two GSE GCs. Additionally, among the neutron-capture elements, this is the chemical space in which Rup106 is the most depleted (see bottom panel of Figure 4). This finding is particularly intriguing, as Cu, the other element where Rup106 is notably depleted, is also produced by massive stars through weak -processes, with a small contribution also by AGB stars (Kobayashi & Nomoto 2009, Kobayashi et al. 2020). Thus, if the lack of weak -process products is responsible for the low Cu derived, it is also expected that light neutron-capture elements would be similarly depleted, as it is observed for Y.
@@
When examining the abundances of the other neutron-capture elements (see Figure 5), the most notable distinctions between various GCs arise in the r-process element Eu. The two GSE GCs exhibit values that are higher compared to in situ GCs. The enhancement in r-process of NGC 362 and NGC 1261 has also been highlighted by Monty et al. (2023b) and by Koch-Hansen et al. (2021) respectively, and it resembles the behaviour of GSE field stars at this metallicity (Nissen & Schuster 2011, Fishlock et al. 2017, Aguado et al. 2021, Matsuno et al. 2021, Naidu et al. 2022b) and that of surviving dwarf galaxies, like the LMC, Sculptor and Fornax (Tolstoy et al. 2009, Letarte et al. 2010, Van der Swaelmen et al. 2013, Lemasle et al. 2014). This might be explained by a joint effect of the different star formation efficiency in the progenitor galaxy and the impact of delayed -process sources (i.e. NSMs), with minimum delay times of the order of - Myr in producing Eu (Cescutti et al. 2015, Siegel et al. 2019, Naidu et al. 2022b, Ou et al. 2024). Once again, Rup106 is underabundant in all neutron-capture elements, thus indicating a different birth environment.
@@
Recently, the [EuII/] ratio has been proposed as a strong discriminant between accreted and in situ GCs due to its effectiveness in tracing different timescales of star formation (Monty et al. 2024, Ou et al. 2024). As shown in the top panel of Figure 6, the two GSE GCs follow the expected behaviour compared to NGC 6218, NGC 6522 and NGC 6626, while we do not find sign of a consistent enhancement in [EuII/] for Rup106. Thus, we recommend caution when using this diagnostic, as it might be effective for individual progenitors like GSE, but not in general for birth environments spanning a wide range of formation and evolutionary properties. Moreover, as demonstrated by the [EuII/BaII] and the [EuII/LaII] ratios (see bottom panel of Figure 6), in situ GCs and Rup106 show signs of higher efficiency in the production of heavy s-process elements with respect to the r-process elements when compared to the GSE GCs. We highlight that accreted systems considered here can be either r-process dominated (GSE) or s-process dominated (Rup106). Moreover, we find a difference of in [EuII/YII] between the three accreted cluster and the three in situ systems. Interestingly, the [EuII/YII] ratio is similar in all accreted systems, regardless of the r-/s- process dominance. This comparison may suggest that the applicability of [EuII/YII] as a tracer of the accreted/in situ origin is more general than [EuII/BaII] and [EuII/LaII]. Thus, this abundance ratio can play an important role in discriminating between in situ and accreted GCs at this metallicity, but, as for Zn, we point out that this diagnostic might lack sensitivity to discriminate between independent accreted progenitors.
@@ 5 Comparison with field stars from the likely progenitor system
In this section, we investigate the compatibility of the abundances derived in accreted GCs with those of field stars associated to their respective progenitors, and in passing we remark once again that homogeneity in this kind of comparison is crucial. In fact, several factors may contribute to a zero-point offset in abundance between our results and those from the literature. Discrepancies in the assumed atmospheric parameters, model atmosphere, solar mixture, linelist, and atomic data (e.g., log gf) can potentially result in variations in the derived abundances.
5.1 NGC 6218, NGC 6522, NGC 6626 and the Milky Way

### v2

@@ 3.2 Deriving the abundances
The determination of chemical abundances of Mg, Si, Ca, TiI, TiII, Cr, Fe, Ni, and Zn was conducted by means of a comparative analysis between the observed equivalent widths (EWs) —measured with the code DAOSPEC (Stetson & Pancino 2008) and using the automatic tool 4DAO (Mucciarelli 2013)— and theoretical line strengths. This analysis was executed using the code GALA (Mucciarelli et al. 2013).
We employed the spectral synthesis using the proprietary code SALVADOR to derive the chemical abundances for the species that have hyperfine or isotopic splitting transitions (ScII, V, Mn, Co, Cu, YII, BaII, LaII, and EuII). SALVADOR runs a -minimisation between the observed line and a grid of suitable synthetic spectra computed on the fly by the code SYNTHE (Kurucz 2005) varying only the abundance of the matching element and keeping the stellar parameters fixed. Synthetic spectra were computed using all the atomic and molecular transitions available in the Kurucz/Castelli333http://wwwuser.oats.inaf.it/castelli/linelists.html database.
We used ATLAS9 (Kurucz 2005) model atmospheres computed assuming plane-parallel geometry, hydrostatic and radiative equilibrium, and local thermodynamic equilibrium for all the chemical elements. We started from an -enhanced model atmosphere ([/Fe] = +0.4 dex) for all the target GCs except for Rup106, for which we used a solar-scaled chemical mixture according to the values of we derived. Examples of fits around absorption lines of elements of interest (i.e. Zn and EuII; see Sects. 4.2 and 4.3) are displayed in Fig. 2. Different synthetic spectra are computed for each star by assuming the atmospheric parameters derived as described in Sect. 3.1, and varying the abundances of Zn and EuII. Finally, we scale the abundance ratios to the solar values using the chemical composition described in Grevesse & Sauval (1998), as ATLAS9 model atmospheres are computed based on this solar mixture (Castelli & Kurucz 2003).
@@ 3.3 Abundance uncertainties
We estimated the uncertainties attributed to EW measurements as the dispersion observed around the mean of individual line measurements divided by the root mean square of the number of lines employed in the analysis.
The internal error on chemical abundances derived with spectral synthesis were quantified using Monte Carlo simulations. To do so, we repeated the analysis on 1000 noisy synthetic spectra obtained adding Poissonian noise in order to mimic the S/N of the real spectra. The uncertainty is estimated as the standard deviation of the abundance distribution of the 1000 noisy synthetic spectra.
Finally, we performed new calculations of the chemical abundances, considering also the uncertainties associated with the atmospheric parameters. This involved varying one stellar parameter at a time while keeping the others fixed. In the end, all the sources of error are added in quadrature. Given that our chemical abundances are expressed as abundance ratios ([X/Fe]), the uncertainties on the iron abundance have been taken into account. We estimated the total uncertainty as the squared sum of these two components. Therefore, the final errors were calculated as:
where is the dispersion around the mean of chemical abundances, is the number of used lines, and are the abundance differences obtained by varying the parameter .
@@ 4 Abundances of target GCs
In this section, we discuss the abundances of -, iron-peak, and neutron-capture elements of our target GCs. We exercise caution when interpreting the abundances of certain chemical elements that may be influenced by the presence of multiple populations within GCs (e.g. O, Na, Mg, and Al; see Bastian & Lardo 2018). Among these species, we focus on Mg, which exhibits minimal dispersion within five out of the six GCs in our sample. This outcome aligns with expectations, particularly considering that most of the targeted systems belong to the low-mass and high-metallicity regime of the MW GC system, where the efficiency of the MgAl burning channel diminishes (Ventura et al. 2013, Dell’Agli et al. 2018, Alvarez Garay et al. 2024). Indeed, we find spread in the distribution of [Mg/Fe] only in the most massive among our target cluster, NGC 6626. As we are only concerned with the pristine chemical composition of the gas that formed the cluster, we derived the mean value of [Mg/Fe] using only first-generation stars. The results obtained from the homogeneous chemical analysis of 66 RGB stars of the sample of six GCs are listed in Tables 2 - 4. The mean abundances for each GC are available in Table 5. The average abundance ratios of the target GCs are illustrated in Figs. 3 - 5.
@@
Fig. 3 shows the mean values of the -elements abundances. These chemical elements are predominantly synthesised in massive stars and subsequently dispersed into the interstellar medium through core-collapse supernovae (CC-SNe, Kobayashi & Nomoto 2009, Romano et al. 2010, Kobayashi et al. 2020) events. Additionally, there is a minor but discernible contribution from Type Ia supernovae (SNe Ia), which is particularly notable for Ca and Ti, as discussed in Kobayashi et al. (2020).
First of all, we find that the two GCs associated to GSE (green filled symbols) show identical abundances, within the uncertainties, in all the analysed -elements. Second, we show that the mean abundances of Mg, Si, and Ca of the three in situ GCs, namely NGC 6218, NGC 6522, and NGC 6626 (red, brown and pink filled symbols, respectively), are fully compatible among each other. For Ti, this consistency is less evident. Third, we find that NGC 362 and NGC 1261 are -depleted compared to the in situ GCs. They display differences in the -element at the level in Mg and Ca when comparing the average abundances and the respective standard deviations (see Appendix A.1 for a comparison with literature). Finally, we derive subsolar abundances in the -elements of Rup106 (blue filled squares), specifically with , which is in good agreement with the literature (Brown et al. 1997, V13; see Appendix A.2 for a comparison with the results of V13). This is, on average, lower than the other five target clusters.
These results guide us as to the degree to which we can use -elements to distinguish the origin of GCs. The observed trend is a good match to our expectations based on the cluster chrono-dynamical associations. Indeed, GSE clusters NGC 1261 and NGC 362 share the same -element abundances, as expected for clusters accreted from the same progenitor, with lower values compared to NGC 6218, NGC 6522, and NGC 6626. Indeed, the in situ GCs consistently show higher abundances, as predicted for systems born in the MW at this metallicity, where the contribution of SNIa to the gas chemical enrichment was not yet significant. Finally, Rup106 has an even lower than the GSE GCs, and can be distinguished from them in this respect; this indicates an even more different birth environment, probably characterised by very low star formation efficiency.
@@ 4.2 Iron-peak elements
Iron-peak elements primarily originate in Type II CC-SNe and hypernovae (HNe), with some contribution also from SNe Ia. Specifically, elements such as Sc, Cu, and Zn are predominantly synthesised by massive stars, while for V and Co the contribution of SNe Ia is not negligible. On the other hand, Cr, Mn, and Ni are primarily produced by SNe Ia (Romano et al. 2010, Kobayashi et al. 2020).
The results shown in Fig. 4 indicate that most of the iron-peak elements do not provide a means to effectively and clearly discriminate between in situ and GSE clusters. Indeed, our target GCs display coherent values in all the chemical spaces with the exception, once again, of Rup106, which is underabundant in ScII, V, Mn, Co, and Ni. In the bottom panel of Fig. 4, we highlight the two elements where Rup106 is most depleted compared to the other systems. Specifically, Cu is the most underabundant among the chemical elements derived in this work ([Cu/Fe] for Rup106). Such low values of [Cu/Fe] can be reproduced in stellar systems with a chemical evolutionary model assuming very inefficient star formation (star formation rate , Mucciarelli et al. 2021b). Thus, the pronounced depletion in all of these elements might indicate that Rup106 was born in an environment where the contribution of massive stars was extremely low.
Among the iron-peak elements, the only one to demonstrate a statistically significant difference between accreted and in situ GCs is Zn: all the in situ GCs are overabundant with respect to GSE GCs and Rup106 at the 1 level. In passing, we note that NGC 6626 shows a significant spread in [Zn/Fe], as already found in Villanova et al. (2017). Zn therefore appears to be a good tracer of the in situ or accreted origin of GCs, but it is not sensitive enough to help distinguish among different accreted progenitors. Such a conclusion was also drawn by Minelli et al. (2021), even though these authors used a sample of GCs at higher metallicity. Indeed, high [Zn/Fe] ratios are expected for in situ GCs, and the rationale behind this expectation lies in the fact that the primary contributors to Zn production are hypernovae, which are linked to stars with masses of around (Romano et al. 2010, Kobayashi et al. 2020). Consequently, galaxies with lower star formation rates should manifest lower [Zn/Fe] ratios due to the reduced impact of massive stars (Yan et al. 2020).
@@ 4.3 Neutron-capture elements
Slow (-) neutron-capture elements are mainly synthesised in asymptotic giant branch (AGB) stars. In particular, light -process elements (e.g. Y) are produced to a large extent by intermediate-mass AGB stars. Conversely, heavy -process elements (e.g. Ba, La) are mainly produced in AGB stars with masses lower than (Kobayashi et al. 2020). Rapid (-) neutron-capture elements (e.g. Eu) are synthesised during a broad spectrum of events, such as special types of CC-SNe (rCCSNe, e.g. magnetorotational SNe, collapsars; Mösta et al. 2018, Siegel et al. 2019, Kobayashi et al. 2020) and neutron-star mergers (NSMs, Lattimer & Schramm 1974).
@@
First of all, we observe that the in situ NGC 6218 is enhanced () in [YII/Fe] compared to the other two in situ GCs and two GSE GCs. Additionally, among the neutron-capture elements, this is the chemical space in which Rup106 is the most depleted (see the bottom panel of Fig. 4). This finding is particularly intriguing, as Cu, the other element where Rup106 is notably depleted, is also produced by massive stars through weak -processes, with a small contribution also by AGB stars (Kobayashi & Nomoto 2009, Kobayashi et al. 2020). Thus, if the lack of weak -process products is responsible for the low derived Cu abundance, light neutron-capture elements are also expected to be similarly depleted, as is observed for Y.
@@
When examining the abundances of the other neutron-capture elements (see Fig. 5), the most notable distinctions between various GCs arise in the r-process element Eu. The two GSE GCs exhibit values that are higher compared to in situ GCs. The enhancement in r-process of NGC 362 and NGC 1261 was also highlighted by Monty et al. (2023b) and by Koch-Hansen et al. (2021), respectively, and resembles a behaviour seen in both GSE field stars at this metallicity (Nissen & Schuster 2011, Fishlock et al. 2017, Aguado et al. 2021, Matsuno et al. 2021, Naidu et al. 2022b) and surviving dwarf galaxies, such as the Large Magellanic Cloud, Sculptor, and Fornax (Tolstoy et al. 2009, Letarte et al. 2010, Van der Swaelmen et al. 2013, Lemasle et al. 2014). This might be explained by a joint effect of the different star formation efficiency in the progenitor galaxy and the impact of delayed -process sources (i.e. NSMs), with minimum delay times of the order of - Myr in producing Eu (Cescutti et al. 2015, Siegel et al. 2019, Naidu et al. 2022b, Ou et al. 2024). Once again, Rup106 is underabundant in all neutron-capture elements, indicating a different birth environment.
@@
Recently, the [EuII/] ratio has been proposed as a strong discriminant between accreted and in situ GCs due to its effectiveness in tracing different timescales of star formation (Monty et al. 2024, Ou et al. 2024). As shown in the top panel of Fig. 6, the observed differences between the two GSE GCs and the in situ GCs are consistent with literature results, while we do not find signs of a consistent enhancement in [EuII/] for Rup106. Thus, we recommend caution when using this diagnostic, as it might be effective for individual progenitors like GSE, but not in general for birth environments spanning a wide range of formation and evolutionary properties. Moreover, as demonstrated by the [EuII/BaII] and the [EuII/LaII] ratios (see the bottom panel of Fig. 6), the in situ GCs and Rup106 show signs of higher efficiency in the production of heavy s-process elements with respect to the r-process elements when compared to the GSE GCs. We highlight that the accreted systems considered here can be either r-process dominated (GSE) or s-process dominated (Rup106). Moreover, we find a difference of in [EuII/YII] between the three accreted clusters and the three in situ systems. Interestingly, the [EuII/YII] ratio is similar in all accreted systems, regardless of the r-/s- process dominance. This comparison may suggest that the applicability of [EuII/YII] as a tracer of the accreted or in situ origin is more general than [EuII/BaII] and [EuII/LaII]. Thus, this abundance ratio can play an important role in discriminating between in situ and accreted GCs at this metallicity, but, as for Zn, we point out that this diagnostic might lack the required sensitivity to discriminate between independent accreted progenitors.
@@ 5 Comparison with field stars from the likely progenitor system
In this Section we investigate the compatibility of the abundances derived for accreted GCs with those derived for the field stars associated with their respective progenitors, and in passing we note once again that homogeneity in this kind of comparison is crucial. Indeed, several factors may contribute to a zero-point offset in abundance between our results and those from the literature. Discrepancies in the assumed atmospheric parameters, model atmosphere, solar mixture, line list, and atomic data (e.g. log gf) can potentially result in variations in the derived abundances.
5.1 NGC 6218, NGC 6522, NGC 6626, and the Milky Way